Is it possible to have a regular polygon each of whose exterior angel is 50 ?
Found 2 solutions by Edwin McCravy, Theo:
Answer by Edwin McCravy(20054)

No, and here's why.  It's because: 

1. There are n exterior angles of any n-sided polygon.

2. All the exterior angles of any polygon have sum 360°.

3. All exterior angles of a REGULAR polygon are congruent, i.e.,
   their measures are all equal.

4.  Therefore each exterior angle of an n-sided polygon is 360°/n.

5.  So therefore if it were possible to have such a regular polygon, 
    then each of its exterior angles would be 50°.

6. That would mean that 360°/n = 50°

7.  When we solve 360°/n = 50° for n we get:
                    360° = 50°n
                360°/50° = n
                  360/50 = n
                    36/5 = n
                   7 1/5 = n

8.  No polygon, regular or not, can have 7 and 1/5th sides.  
    The number of sides is always a counting number, a positive
    whole number or integer.

Edwin


Answer by Theo(13342)

i don't think so.
i believe the number of sides of a regular polygon has to be an integer.
the number of sides is equal to 360 divided by the exterior angle.
360/50 is equal to 7.2 which is not an integer.